Show that if events A and B are independent, then the following pairs of events are also independent. (a) A and B¯ (b) A¯ and B¯.


1
Expert's answer
2021-03-01T07:16:52-0500

P(A∩B) = P(A) P(B) by the definition of independence


= P(A) (1-P(B')) since P(B) = 1- P(B')


= P(A) - P(A) P(B') 


So,

 

(1) P(A) P(B') = P(A) - P(A∩B) 

 

Since A∩B' = A - A∩B and A∩B ⊂ A, 

 

(2) P(A∩B') = P(A) - P(A∩B)

 

From (1) and (2), P(A∩B') = P(A) P(B'), so A and B' are independent

 

This argument shows that if two events are independent, then each event is independent of the complement of the other. Therefore A' and B' are also independent events.